# How Visual Representations Map to Language Feature Space in Multimodal LLMs

### v1

@@ Abstract
Effective multimodal reasoning depends on the alignment of visual and linguistic representations, yet the mechanisms by which vision-language models (VLMs) achieve this alignment remain poorly understood. We introduce a methodological framework that deliberately maintains a frozen large language model (LLM) and a frozen vision transformer (ViT), connected solely by training a linear adapter during visual instruction tuning. This design is fundamental to our approach: by keeping the language model frozen, we ensure it maintains its original language representations without adaptation to visual data. Consequently, the linear adapter must map visual features directly into the LLM’s existing representational space rather than allowing the language model to develop specialized visual understanding through fine-tuning. Our experimental design uniquely enables the use of pre-trained sparse autoencoders (SAEs) of the LLM as analytical probes. These SAEs remain perfectly aligned with the unchanged language model and serve as a snapshot of the learned language feature-representations. Through systematic analysis of SAE reconstruction error, sparsity patterns, and feature SAE descriptions, we reveal the layer-wise progression through which visual representations gradually align with language feature representations, converging in middle-to-later layers. This suggests a fundamental misalignment between ViT outputs and early LLM layers, raising important questions about whether current adapter-based architectures optimally facilitate cross-modal representation learning.  [^1] Trained weights and code available at https://github.com/cvenhoff/vlm-mapping
@@ 1 Introduction
To explore this, we apply mechanistic interpretability tools [4, 21] to analyze VLM internal representations beyond behavioral outputs. Central to our method are sparse autoencoders (SAEs) [1], which disentangle neural representations into interpretable features. We use SAEs to study cross-modal mapping by connecting a frozen vision transformer (CLIP ViT-L14 [19]) and a frozen LLM (Gemma-2-2b-it [24]) through a trainable linear projector. By keeping both the ViT and LLM frozen, we enforce a setting where the projector must align visual inputs to the LLM’s fixed feature space, without altering the language model’s internal structure.
@@
Our results unveil the middle-to-late layers as the convergence location for the cross-modal mapping by finding that the semantic alignment of SAE feature descriptions and the visual input steeply increases in these layers, while the SAE reconstruction error and sparsity decrease. This reveals that visual representations gradually become represented by precise, sparse language model features, indicating successful cross-modal mapping from vision to language model feature space. Crucially, this goes beyond simple out-of-distribution detection as it reveals the specific language model features that the adapter uses to represent visual information in the middle—to-late layers.
@@ 2.1 Multimodal Mechanistic Interpretability
Recent multimodal interpretability research has explored vision-language models through various approaches. [2] and [7] analyzed CLIP, while [20] examined multimodal neurons in VLMs. [10] investigated VLM responses to hallucinated versus real objects. For mechanistic interpretation of MLLMs, [18] used Causal Mediation Analysis on BLIP, [16] projected visual representations onto language vocabulary, while [26] probed for the emergence of visual representations in the language model backbone. However, these approaches, particularly [16]’s logit lens methodology, focus primarily on the embedding or unembedding space without investigating how information propagates through inner representations—a limitation our work addresses using sparse autoencoders to analyze feature representation across all network layers.
@@ 2.2 Dictionary Learning and Sparse Autoencoders
Sparse Dictionary Learning (SDL) methods [8, 5] decompose complex models into fundamental components, identifying more directions than available dimensions—ideal for neural networks exhibiting superposition [5]. Sparse Autoencoders (SAEs) [1], a type of SDL, enable comprehensive analysis across all network layers through enforced activation sparsity. To avoid the computational cost of training our own SAEs, we use Gemma Scope [12], pre-trained SAEs for Gemma2-2b [23].
@@ 3 Training the Cross-Modal Adapter
In this section, we detail the methodology used to train and evaluate our vision language model (VLM). Our approach follows a controlled experimental design by keeping both the vision transformer and the language model frozen during training, ensuring that the learned adapter is solely responsible for mapping visual representations into the existing language feature space.
@@ 3.1 Training Setup
We employ a LLaVA-style [14] adapter training setup, where a vision transformer and a LLM are connected via a learnable linear projection layer. However, unlike conventional LLaVA-style VLM training, we freeze both the vision transformer (ViT) and the large language model (LLM) throughout all training stages. This setup enables precise interpretability of the cross-modal mapping process.
@@ 4.1 SAE-Based Analysis Framework
where $||\cdot||_{0}$ counts the number of nonzero elements and $d_{sae}$ is the dictionary size of the SAE. If sparsity levels for visual representations are comparable those for language representations, this suggests that the visual representations have mapped to the language model features.
@@ 5 Conclusion
This work presented a methodological framework for analyzing the mapping of visual features to language representations in multimodal LLMs. By connecting a vision encoder to a frozen language model through a trainable projector, we created a controlled environment where the connector must learn to translate visual information into the language model’s existing semantic space. By leveraging pre-trained sparse autoencoders for the frozen LLM, we tracked how visual information flows through the language model, identifying where vision tokens become aligned with language representations. Our reconstruction error analysis and semantic feature investigation revealed how the connector bridges modalities, demonstrating when and how visual information is mapped to language semantics.

### v2

@@ Abstract
Effective multimodal reasoning depends on the alignment of visual and linguistic representations, yet the mechanisms by which vision-language models (VLMs) achieve this alignment remain poorly understood. Following the LiMBeR framework, we deliberately maintain a frozen large language model (LLM) and a frozen vision transformer (ViT), connected solely by training a linear adapter during visual instruction tuning. By keeping the language model frozen, we ensure it maintains its original language representations without adaptation to visual data. Consequently, the linear adapter must map visual features directly into the LLM’s existing representational space rather than allowing the language model to develop specialized visual understanding through fine-tuning. Our experimental design uniquely enables the use of pre-trained sparse autoencoders (SAEs) of the LLM as analytical probes. These SAEs remain perfectly aligned with the unchanged language model and serve as a snapshot of the learned language feature-representations. Through systematic analysis of SAE reconstruction error, sparsity patterns, and feature SAE descriptions, we reveal the layer-wise progression through which visual representations gradually align with language feature representations, converging in middle-to-later layers. This suggests a fundamental misalignment between ViT outputs and early LLM layers, raising important questions about whether current adapter-based architectures optimally facilitate cross-modal representation learning.  [^1] Trained weights and code available at https://github.com/cvenhoff/vlm-mapping
@@ 1 Introduction
To explore this, we apply mechanistic interpretability tools [4, 24] to analyze VLM internal representations beyond behavioral outputs. Central to our method are sparse autoencoders (SAEs) [1], which disentangle neural representations into interpretable features. We build on the LiMBeR framework that demonstrates effective feature mapping from image to language space [17]), by connecting a frozen vision transformer (CLIP ViT-L14 [22]) and a frozen LLM (Gemma-2-2b-it [27]) via a trainable linear adapter. By keeping both the ViT and LLM frozen, we enforce a setting where the adapter must align visual inputs to the LLM’s fixed feature space, without altering the language model’s internal structure.
@@
Our results unveil the middle-to-late layers as the convergence location for the cross-modal mapping by finding that the semantic alignment of SAE feature descriptions and the visual input steeply increases in these layers, while the SAE reconstruction error and sparsity decrease. This reveals that visual representations gradually become represented by precise, sparse language model features, indicating successful cross-modal mapping from vision to language model feature space. Crucially, this goes beyond simple out-of-distribution detection as it reveals the specific language model features that the adapter uses to represent visual information in the middle-to-late layers.
@@ 2.1 Multimodal Mechanistic Interpretability
Recent interpretability research has explored vision-language models through various approaches. [2] and [8] analyzed CLIP, while [23] examined multimodal neurons in VLMs. [11] investigated VLM responses to hallucinated versus real objects. [20] used Causal Mediation Analysis on BLIP, [18] projected visual representations onto language vocabulary, while [29] probed for the emergence of visual representations in the language model backbone.
@@ 2.2 Dictionary Learning and Sparse Autoencoders
Sparse Dictionary Learning (SDL) methods [9, 6] decompose model activations into interpretable directions. Sparse Autoencoders (SAEs) [1], a type of SDL, enable comprehensive analysis across all network layers through enforced activation sparsity. To avoid the computational cost of training our own SAEs, we use Gemma Scope [13], pre-trained SAEs for Gemma2-2b [26].
2.3 Linear Mapping Approaches
LiMBeR [17] introduced the setup of mapping frozen vision encoder outputs into a frozen LLM’s input space via a trained linear projection, enabling image captioning without tuning either backbone. [21] improves on this with an optimal transport loss aligning visual and text representations to word embeddings. [5] trains only a gated nonlinear adapter for efficient fusion.
@@ 3 Training the Cross-Modal Adapter
In this section, we detail the methodology used to train and evaluate our vision language model (VLM). Following the LiMBeR framework [17] we keep both the vision transformer and the language model frozen during training, ensuring that the learned adapter is solely responsible for mapping visual representations into the existing language feature space.
@@ 3.1 Training Setup
We employ a LLaVA-style [15] adapter training setup, where a vision transformer and a LLM are connected via a learnable linear projection layer. As discussed, we freeze both the vision transformer (ViT) and the large language model (LLM) throughout all training stages. This setup enables precise interpretability of the cross-modal mapping process.
@@ 4.1 SAE-Based Analysis Framework
where $||\cdot||_{0}$ counts the number of nonzero elements and $d_{sae}$ is the dictionary size of the SAE. If sparsity levels for visual representations are comparable to those of language representations, then this suggests that the visual representations have mapped to the language model features.
@@ 5 Conclusion
This work presented a methodological framework for analyzing the mapping of visual features to language representations in multimodal LLMs. Following the LiMBeR framework [17], we connect a frozen vision encoder to a frozen language model through a trainable projector, to force the adapter to learn to translate visual information into the language model’s existing semantic space. Then we leverage pre-trained sparse autoencoders for the frozen LLM to track how visual information flows through the language model, identifying where vision tokens become aligned with language features. Our reconstruction error analysis and semantic feature investigation revealed how the connector bridges modalities, demonstrating when and how visual information is mapped to language semantics.